# Motion Blender Gaussian Splatting for Dynamic Reconstruction

### v1

@@ 3.3 Weight Painting Function
To estimate the influence of each graph link on a Gaussian, we represent the weight painting function as: $\mathcal{W}(\mathbf{x}_{0},X_{\mathcal{L},0})=\mathrm{softmax}(\{K(\mathbf{x}_{0},X_{\mathcal{L},0,i})\mid\forall i\in[1,L]\})$, where $\mathbf{x}_{0}$ denotes the initial position of a Gaussian, and $X_{\mathcal{L},0,i}\in\mathbb{R}^{6}$ represents the initial positions of the start and end joints of the $i$-th link. The kernel function $K_{\psi}:\mathbb{R}^{3}\times\mathbb{R}^{6}\to\mathbb{R}$, measures the affinity between a Gaussian and a link. We define the kernel as:
@@
where $\gamma$ is a learnable parameter that controls the kernel radius. By learning distinct $\gamma$ at each link, the motion graph’s influence regions collectively span the object’s surface. At $t=0$, 3D Gaussians reconstruct the objects’ initial geometry. The static reconstruction is bonded to the dynamic motion graph through weight painting. The painted weights are subsequently used to propagate motions for $t>0$ to achieve dynamic reconstruction.
3.4 Optimization Details
Figure 5: Novel Poses from Motion Graph Editing. The first row shows the observed scene rendered from reconstructed Gaussians. By editing motion graphs and propagating changes to the Gaussians, we create novel, unseen poses, as rendered in the second row.
To reconstruct dynamic scenes from videos, we follow common practices [44, 29] and select the frame where objects occupy the largest image area as the canonical frame instead of using the first frame. For simplicity, we still use $t=0$ to denote canonical frame. We extract instance segmentation masks using Grounding SAM2 [26, 19, 37, 14, 38, 36] and 2D human skeletons using SAPIENS [18]. These annotations are used to initialize the motion graph for each instance at the canonical frame. Specifically, the motion graph is parameterized by $(\theta_{t})_{t=0}^{T-1}$ (time-varying parameters) and $\phi$ (time-independent parameters). We create the graph for each instance at the canonical frame with $(\theta_{0},\phi)$ and propagate the parameters across time by setting $\theta_{t}=\theta_{0},\forall t\in[0,T-1]$. We illustrate the motion graph initialization procedure for the canonical frame in Fig. 6. The 3D Gaussians are initialized for every instance and background using procedure from prior work [44]. At each iteration, we transform the 3D Gaussians from the canonical frame $t=0$ to a target frame $t$ using Eq. 2, and minimize the rendering loss against the corresponding video frame.
Figure 6: Motion Graph Initialization. At $t=0$, we lift 2D skeletons and instance masks to 3D point clouds using depth images. For kinematic structures such as the human body, we leverage prior knowledge of joint connections to initialize a kinematic tree with standard joint rotations and link lengths. This tree is then fitted to the point cloud skeleton by minimizing the average point-to-link distance via gradient descent. For deformable graphs, we sample points from the instance point cloud using farthest point sampling [6] and connect adjacent points to form the graph.
The $i$-th row of $M$ represents the one-hot encoding of the instance index for each Gaussian. We then splat $M$ into a 2D instance mask of dimension $\mathbb{R}^{H\times W\times I}$ and minimize the L1 distance between this rendered mask and the instance mask predicted by Grounding SAM2. In doing so, 3D positions of Gaussians are grouped to faithfully represent the geometry of each instance, as illustrated in Fig. 10.
2D Keypoints Regularization. We project the 3D joints of the human kinematic tree onto the image plane and minimize the L1 distance between the projected joint positions and the 2D human keypoints predicted by SAPIENS.
@@ 4 Experiment
Sec. 4.1 introduces our experimental settings. In Sec. 4.2, we compare our results with SoTA Gaussian splatting methods on real-world datasets. We then show two applications of our method: creating novel object poses and synthesizing robot demonstrations. Finally, we conduct ablation studies in Sec. 4.4 and discuss the limitations in Sec. 4.5.
@@ 4.1 Settings
Datasets. We evaluate our method on two real-world datasets: the iPhone dataset [8] from [44] and the HyperNeRF [32] vrig dataset. The iPhone dataset consists of 12 scenes, including 5 multi-camera (MV) and 7 single-camera (SV) scenes, each captured at $960\times 720$ resolution. The MV scenes are captured using one hand-held moving camera and two fixed cameras at novel angles, where the hand-held video is used for training and the fixed cameras for quantitative evaluation. The SV scenes, captured solely with a hand-held camera, are used for qualitative visualization. The HyperNeRF vrig dataset consists of 4 scenes, each captured at $960\times 540$ resolution using two cameras mounted on a rig kit with strong camera motions. Models are trained on a subset of frames and evaluated on the remaining frames.
Baselines. On the iPhone dataset, we compare against Shape-of-Motion [44]. For HyperNeRF, we compare against 4DGaussians [46]. Both represent the SoTA on their respective datasets. Following Shape-of-Motion, we use their proposed 2D track loss in addition to L1 rgb loss.
@@ 4.2 Results
Figure 7: Visualization of novel view rendering on iPhone dataset compared with other methods. Regions in green are excluded from evaluation, as these pixels are never seen in training videos. LPIPS on bottom left with best colored red.
Table 1: Quantitative results for novel view rendering on iPhone dataset. LPIPS more accurately reflects perceptual quality.
@@
Figure 8: Visualization of results in HyperNerf dataset compared with other methods. LPIPS on bottom left with best colored red.
Table 2: Quantitative results on HyperNerf dataset. Our method performs competitively, closely matching SoTAs in LPIPS.
@@
Tab. 1 presents our quantitative results on the iPhone benchmark, where our method outperforms the state-of-the-art Shape-of-Motion by 0.02 LPIPS. Fig. 7 provides a qualitative comparison on the five MV scenes with 4DGaussians [46] and Shape-of-Motion [44], demonstrating that our method renders sharper, more complete, and perceptually higher-quality novel views. While 4DGaussians achieves comparable PSNR and SSIM scores, its rendering quality is notably inferior. Fig. 4 visualizes the learned motion graphs and weight paintings. Training on the iPhone dataset converges within 10 to 30 hours on a 40GB A100 GPU, depending on scene complexity. For the teddy bear scene with 2M Gaussians—using a deformable graph for the bear and a kinematic tree for the human—our method achieves a rendering speed of 18 FPS, and 25 FPS if motion graphs for each frame are pre-computed and cached. On smaller scenes, such as the chicken toy in HyperNeRF with 300K Gaussians, the rendering speed increases to 32 FPS without caching and 46 FPS with caching.
Tab. 2 presents our quantitative results on the HyperNeRF vrig benchmark, with qualitative comparison across four scenes in Fig. 8. While 4DGaussians, the SoTA on HyperNeRF, achieves higher PSNR and SSIM, our method achieves a comparable or better rendering quality and LPIPS on the chicken, 3D printer, and broom scenes (rows 2 to 4 in Fig. 8). Notably, 4DGaussians achieves a better LPIPS score on the 3D printer scene, but our method produces clearer results, accurately rendering the text on the printer motor, unlike the blurry output from 4DGaussians. However, on the peel-banana scene, our method produces lower-quality results than 4DGaussian. We further discuss our limitations in Sec. 4.5.
@@ 4.3 Applications
Figure 9: Robot Demonstration from Human Videos. We reconstruct Gaussians from human and robot videos, with the robot’s pre-defined kinematic chain as the motion graph. We then remove Gaussians of the human and add robot’s to workspace. Next, we animate the robot motion graph using inverse kinematics derived from the hand pose. This renders videos of a robot arm mimicking human actions.
@@
Fig. 5 shows that our method extends beyond reconstructing observed motions, enabling the creation of novel object motions while maintaining high rendering quality. Recent advances in controllable image and video editing predominantly rely on large diffusion models [2, 41, 11]. In contrast, our method introduces a novel approach by operating directly in 3D space, enabling precise control over object poses without the need for neural networks.
Figure 10: Visualization of rendered scene (left) and underlying Gaussian positions as point cloud (right). The point cloud is colored to distinguish individual instances and background.
Robot Demonstration Synthesis. Imitation learning enables robots to perform complex tasks by learning from robot demonstrations. Unlike human videos, collecting robot demonstrations requires labor-intensive teleoperation. In Fig. 9, we demonstrate a prototype of our method for synthesizing robot demonstrations from human videos.
@@
We reconstruct scenes from human and robot videos independently. For robot videos, we use the pre-defined robot kinematic chain as the motion graph, available from the robot system runtime. We then remove the human Gaussians and add robot Gaussians to the workspace. Keyframes are manually selected, with hand positions mapped to robot gripper poses using our visual editor. Gripper poses between keyframes are interpolated, and inverse kinematics converts them into joint angles of the robot kinematic chain. The chain is then used as the robot motion graph for view rendering. This generates a video where the robot replaces the human, completing the task as demonstrated. While artifacts and unrealistic interactions still present, this demonstrates the potential of our approach for robotic applications.
4.4 Ablation Studies
@@
Figure 11: Relationship between novel view rendering quality and motion graph size.
@@
Figure 12: Visualization of rendered images and motion graphs with and without 2D human keypoint regularization (left) and canonical frame regularization (right).
@@
Motion Graph Size. Fig. 11 illustrates the relationship between novel view rendering quality and the number of joints in a motion graph. We observe that the best rendering quality for the teddy scene is achieved with a motion graph of 200 joints. However, increasing the number of joints to 1000 leads to a noticeable drop in quality, suggesting the existence of an optimal graph size. This finding highlights the potential of learning adaptive graph structures that dynamically adjust to different objects.
@@
2D Keypoints Regularization. Fig. 12 compares rendered images and motion graphs with and without 2D human keypoint regularization. We observe that enforcing consistency between the projected 3D kinematic tree and 2D human keypoints (detected by SAPIENS [18]) improves performance, reducing LPIPS by 0.01 to 0.02. More importantly, this regularization produces cleaner and more meaningful motion graphs that align well with the human 3D structure. Without 2D keypoint regularization, the motion graph tends to drift or produce suboptimal structures. For example, if two fingers move together, the kinematic tree may incorrectly control both with a single finger link, leaving other finger links static. This highlights the importance of 2D keypoint regularization in learning accurate and interpretable 3D motion graphs.
@@
Canonical Frame Regularization. The right side of Fig. 12 compares rendered images and motion graphs with and without canonical frame regularization, which minimizes the distance between the learned motion graph’s joint positions and their initial positions at the canonical frame. The joints and links of the motion graph are color-coded based on their distance to the Gaussians on the object surface. With canonical frame regularization, the motion graph remains closely aligned with the object surface. Without it, however, the graph joints become spiky (indicated by red-colored joints) and drift outward, detaching from the object.
4.5 Limitations
Fig. 13 illustrates examples of imperfectly learned motion graphs and a failure case in motion editing. Despite using 2D human keypoint regularization, learning precise and complex kinematic trees for small regions remains challenging. For instance, in the left image, the hand—which contains intricate kinematic structures—occupies a small region (less than $80\times 80$ pixels in a $720\times 920$ image), making it difficult to capture fine details. Additionally, the learning of deformable graphs heavily depends on the quality of the canonical frame’s point cloud. Noisy point clouds can lead to suboptimal initializations, such as joints placed outside the object or incomplete coverage of certain object parts, as shown in the middle example. Further, the right example demonstrates a failure case in motion editing, where attempting to turn the cat’s head introduces non-negligible visual artifacts. Unlike meshes, Gaussians lack an explicit surface representation, allowing them to deform arbitrarily when motion graphs are applied. This can result in unrealistic deformations, especially for motions not present in the training videos.
Figure 13: Visualization of imperfect learned motion graphs (left, middle) and a failure case in motion editing (right).
In addition to the limitations in Fig. 13, our method produces inferior rendering quality on certain HyperNeRF scenes. HyperNerf scenes are captured with strong rotating camera motions, where objects frequently enter and exit the scene, which is deemed unrealistic by [8]. 4DGaussians addresses this challenge by employing a deformation field network, allowing all Gaussians to move freely. In contrast, our method does not use neural networks and learns motion graph parameters directly at each frame, making it less adaptable to rapid camera motions. A promising direction is to integrate a deformation network into the motion graph framework, predicting motion graph parameters instead of learning them as free variables, enhancing adaptability while preserving interpretability.
5 Conclusion
We introduced Motion Blender Gaussian Splatting (MB-GS), a framework for dynamic scene reconstruction using explicit and sparse motion graphs. MB-GS achieves state-of-the-art performance on the iPhone dataset and competitive results on HyperNeRF. We demonstrate its applications in novel pose animation and synthesizing robot demonstrations from human videos. We hope our work can be useful for future research in dynamic scene reconstruction, robotics, and animation.

### v2

@@ 2.3 Weight Painting Function
To estimate the influence of each graph link on a Gaussian, the weight painting function is given as: $\mathcal{W}(\mathbf{x}_{0},X_{\mathcal{L},0})=\mathrm{softmax}(\{K(\mathbf{x}_{0},X_{\mathcal{L},0,i})\mid\forall i\in[1,|\mathcal{L}|]\})$, where $\mathbf{x}_{0}$ is the initial position of a Gaussian, and $X_{\mathcal{L},0,i}\in\mathbb{R}^{6}$ are the initial positions of start and end joints of the $i$-th link. The kernel function $K:\mathbb{R}^{3}\times\mathbb{R}^{6}\to\mathbb{R}$ measures the affinity between a Gaussian and a link and is defined as:
@@
where $\gamma$ is a learnable parameter that controls the kernel radius, and $\text{dist}(a,b)$ is the distance between point $a$ to line segment $b$. By learning distinct $\gamma$ at each link, the motion graph’s influence regions collectively span the object’s surface. At $t=0$, 3D Gaussians reconstruct objects’ initial geometry. The static reconstruction is bonded to the dynamic motion graph through weight painting. The painted weights are subsequently used to propagate motions for $t>0$, for dynamic reconstruction.
Figure 5: Novel Poses from Motion Graph Manipulation. The first row shows images of scenes reconstructed from observations. By applying new control actions on the motion graphs and propagating changes to the Gaussians, novel unseen poses are imagined and rendered in the second row.
(a) (a) Motion Graph Initialization
(b) (b) Instance-level Reconstruction
Figure 6: We initialize motion graphs at the canonical frame ($t=0$) using instance segmentation masks from Grounding SAM2 [23, 24] and 2D human skeletons estimated by SAPIENS [25]. Furthermore, our framework enables per-instance reconstruction, where Gaussians are explicitly grouped to maintain accurate instance geometry — a capability previously unexplored in existing literature of dynamic Gaussian splatting. More optimization details are provided in Appendix A.2.
@@ 3.1 Settings
Datasets. We evaluate our method on two real-world datasets: the highly challenging iPhone dataset [15] from [1] and the HyperNeRF [16] vrig dataset. The iPhone dataset consists of 12 scenes, including 5 multi-camera (MV) and 7 single-camera (SV) scenes, each captured at $960\times 720$ resolution. The MV scenes are captured using one hand-held moving camera and two fixed cameras at novel angles, where the hand-held video is used for training and the fixed cameras for quantitative evaluation. The SV scenes, captured solely with a hand-held camera, are used for qualitative visualization. The HyperNeRF vrig dataset consists of 4 scenes, each captured at $960\times 540$ resolution using two cameras mounted on a rig kit with strong camera motions. Models are trained on a subset of frames and evaluated on the remaining frames.
Baselines. On the iPhone dataset, we compare against Shape-of-Motion [1]. For HyperNeRF, we compare against 4DGaussians [3]. Shape-of-Motion and 4DGaussians are the state of the art on their respective datasets. We use the 2D track loss proposed in [1], in addition to L1 RGB loss.
@@ 3.2 Results
Table 1: Novel view rendering on the highly challenging iPhone dataset [1]. LPIPS more accurately reflects perceptual quality.
@@
Table 2: HyperNerf [16]. Our method performs competitively, closely matching SoTA in the key LPIPS metric.
@@
Tab. 1 presents our quantitative results on the iPhone benchmark, where our method outperforms the state-of-the-art Shape-of-Motion by 0.02 LPIPS. Fig. A4 provides a qualitative comparison on the five MV scenes with 4DGaussians [3] and Shape-of-Motion [1], demonstrating that our method renders sharper, more complete, and perceptually higher-quality novel views. While 4DGaussians achieves comparable PSNR and SSIM scores, its rendering quality is notably inferior. Fig. 4 visualizes the learned motion graphs and weight paintings. Training on the iPhone dataset converges within 10 to 30 hours on a 40GB A100 GPU, depending on scene complexity. For the teddy bear scene with 2M Gaussians—using a deformable graph for the bear and a kinematic tree for the human—our method achieves a rendering speed of 18 FPS, and 25 FPS if motion graphs for each frame are pre-computed and cached. On smaller scenes, such as the chicken toy in HyperNeRF with 300K Gaussians, the rendering speed increases to 32 FPS without caching and 46 FPS with caching.
Tab. 2 presents our quantitative results on the HyperNeRF vrig benchmark, with qualitative comparison in Fig. A5. While 4DGaussians, the SoTA on HyperNeRF, achieves higher PSNR and SSIM, our method achieves a comparable or better rendering quality and LPIPS on the chicken, 3D printer, and broom scenes (rows 2 to 4 in Fig. A5). Notably, 4DGaussians achieves a better score on the 3D printer scene, but our method produces clearer results, accurately rendering the text on the printer motor, unlike the blurry output from 4DGaussians. However, on the peel-banana scene, our method produces lower-quality results than 4DGaussian. We further discuss our limitations in Sec. 5.
@@ 3.3 Applications
Figure 9: Robot Demonstration from Human Videos. We reconstruct Gaussians from human and robot videos, with robot’s pre-defined kinematic chain as motion graph. We then remove human’s Gaussians and add robot’s to workspace. Next, we animate the robot motion graph using inverse kinematics derived from hand poses. This renders videos of a robot arm mimicking human actions.
Figure 10: Robotic Action Prediction via Visual Planning. We reconstruct object Gaussians (cloth, microwave, rope) from human videos, simulate potential motion graph trajectories, and render the resulting object images. The optimal trajectory is selected by maximizing PSNR between rendered images and the target goal image. Key joint trajectories (yellow) are mapped to end-effector positions. Videos of our robot experiments can be found at mlzxy.github.io/motion-blender-gs.
@@
Fig. 5 shows that our method extends beyond reconstructing observed motions, enabling the creation of novel imagined object motions while maintaining high rendering quality. Recent advances in controllable image and video editing predominantly rely on large diffusion models [30, 31, 32]. In contrast, our method introduces a novel approach by operating directly in the 3D space, enabling fast and precise control over object poses without the need for neural networks.
@@
Robot Demonstration Synthesis. Imitation learning enables robots to perform complex tasks by learning from robot demonstrations [33]. Unlike human videos, collecting robot demonstrations requires labor-intensive teleoperation [34]. In Fig. 9, we demonstrate a prototype of our method for synthesizing robot demonstrations from human videos. We reconstruct scenes from human and robot videos independently. The human videos feature an operator performing pick-and-place and cloth folding tasks. The robot videos feature a Kuka iiwa robot that moves randomly. We use the robot’s kinematic chain as the motion graph. We then remove the human Gaussians and add the robot Gaussians to the workspace. Keyframes are manually selected, with hand positions adjusted and mapped to robot gripper poses using our visual editor. Gripper poses between keyframes are interpolated, and converted into joint angles of the robot’s kinematic chain using inverse kinematics. The chain is then used as the robot motion graph for view rendering. This generates a video where the robot replaces the human, completing the task as demonstrated. While some artifacts and unrealistic interactions still exist, this demonstrates the potential of our method for robotic data synthesis. The videos are obtained with minimal effort, and can be used for learning vision-based policies.
Robot Visual Planning. Predicting action outcomes is a fundamental human capability that enables complex manipulation skills, such as manipulating objects until a certain goal is reached [35]. While recent work has explored video generation models to incorporate this ability for robotic manipulation [36, 37], these approaches often require complicated pipelines. In Fig. 10, we present a simple visual planning prototype for goal-conditioned manipulation of deformable (cloth and rope) and articulated (microwave door) objects. We first reconstruct object-specific Gaussians from human videos. For each test scene, the motion graph is learned by minimizing the L1 rendering loss from our reconstructed Gaussians, while keeping Gaussians and weights fixed. This process typically converges within a minute. We then simulate various graph trajectories and render the resulting object images. The optimal trajectory is selected by maximizing PSNR with the provided target goal image. In summary, our method learns a dynamic and photorealistic model of the manipulated object on the fly, using a very small number of frames, and performs planning in simulation with the learned model to select optimal actions. The reward function is simply provided in the form of a goal image. This prototype demonstrates the potential of our method for developing more data-efficient model-based robot learning solutions. More experiment details are provided in Appendix A.3.
4 Final Remarks
We conducted ablation studies analyzing motion graph sizes and regularization strategies. Ablation studies, related work discussions, more visualizations and optimization details are included in Appendix A due to space limit. We discuss the limitations and possible future directions in Sec. 5.
Appendix A Appendix
A.1 Related Works
Figure A1: Demonstration of our KUKA robot arm performing cloth folding (left), microwave door adjusting (middle) and rope bending (right), based on planning with our reconstructed motion graphs, without requiring any robot teleoperation data.
Robotics Applications of Dynamic Gaussian Splatting. Recent work has increasingly adopted Gaussian splatting for robotic perception and manipulation. Popular approaches like GaussianGrasper [58] and SplatMover [59] focus on static reconstructions of rigid objects for digital twin applications. For dynamic scenes, ManiGaussian [60] introduces a Gaussian-based world model for manipulation tasks, while EmbodiedGaussians [61] incorporates particle physics for interaction modeling. PhysTwin [62] further advances deformable object reconstruction through mass-spring physics. Despite their impressive results, our work differs in two key aspects: (1) We propose a sparse motion graph representation, while all of these existing work still use dense per-Gaussian motions. (2) Our framework generalizes beyond constrained robotic workspace settings to complex arbitrary scene. We validate our approach on challenging vision benchmarks including iPhone [15] and HyperNeRF [16], demonstrating broader applicability.
A.2 Optimization Details
To reconstruct dynamic scenes from videos, we follow common practices [1, 21] and select the frame where objects occupy the largest image area or the frame with most valid 2D tracks as the canonical frame instead of using the first frame. For simplicity, we still use $t=0$ to denote canonical frame. We extract instance segmentation masks using Grounding SAM2 [63, 64, 65, 66, 23, 24] and 2D human skeletons using SAPIENS [25]. These annotations are used to initialize the motion graph for each instance at the canonical frame. Specifically, the motion graph is parameterized by $(\theta_{t})_{t=0}^{T-1}$ (time-varying parameters) and $\phi$ (time-independent parameters). We create the graph for each instance at the canonical frame with $(\theta_{0},\phi)$ and propagate the parameters across time by setting $\theta_{t}=\theta_{0},\forall t\in[0,T-1]$. We illustrate the motion graph initialization procedure for the canonical frame in Fig. 6 (a). The 3D Gaussians are initialized for every instance and background using procedure from prior work [1]. At each iteration, we transform the 3D Gaussians from the canonical frame $t=0$ to a target frame $t$ using Eq. 2, and minimize the rendering loss against the corresponding video frame.
Motion Graph Initialization. At $t=0$, we lift 2D skeletons and instance masks to 3D point clouds using depth images. For kinematic structures such as the human body, we leverage prior knowledge of joint connections to initialize a kinematic tree with standard joint rotations and link lengths. This tree is then fitted to the point cloud skeleton by minimizing the average point-to-link distance via gradient descent. For deformable graphs, we sample points from the instance point cloud using farthest point sampling [22] and connect adjacent points to form the graph.
The $i$-th row of $M$ represents the one-hot encoding of the instance index for each Gaussian. We then splat $M$ into a 2D instance mask of dimension $\mathbb{R}^{H\times W\times I}$ and minimize the L1 distance between this rendered mask and the instance mask predicted by Grounding SAM2. In doing so, 3D positions of dynamic Gaussians are grouped to always faithfully represent the geometry of each instance, as illustrated in Fig. 6 (b). To the best of our knowledge, our work is the first to study simultaneous per-instance reconstruction in dynamic Gaussian splatting.
2D Keypoints Regularization. We project the 3D joints of the human kinematic tree onto the image plane and minimize the L1 distance between the projected joint positions and the 2D human keypoints predicted by SAPIENS. For our robotic experiments with kinematic trees on non-humanoid objects, see implementation details in Appendix A.3.
A.3 More Details of Real Robot Experiments
We evaluate our method with three real robot manipulation tasks, cloth folding, microwave door adjusting, and rope bending as shown in Fig. 10. We reconstruct the green cloth using a deformable graph, while applying kinematic trees for the microwave and rope. The canonical frame is set to the first frame for all tasks. We use SAM2 for object mask extraction, where we initialize the mask by clicking the object’s central region in the first frame, then propagate it across the entire video. We capture the green cloth scene using three D415 RGB-D cameras, while the microwave and rope scenes are each recorded with a single D415 RGB-D camera. All the three dynamic objects are reconstructed from human demonstration videos.
For the green cloth, we initialize the motion graph at $t=0$ using a rectangular mesh generated via NanoMesh [67] (Fig. 10, first column). For the microwave and rope—which lack standardized keypoint structures—we manually define their kinematic trees. The microwave is represented as a 2-link graph, with 3D keypoints annotated in the first ($t=0$) and final ($t=T-1$) frames. For the rope, we annotate ten 2D keypoints at the first frame, and track them across frames using CoTracker [68]. The rope’s kinematic tree is initialized by projecting the ten manually annotated 2D keypoints at the first frame into 3D space using depth information, with the central keypoint serving as the root. During optimization, we apply 2D keypoint regularization to the rope but omit it for the cloth and microwave.
Given a test scene, we reconstruct its motion graph by optimizing the motion graph parameters of our learned dynamic Gaussian model. The optimization minimizes the L1 distance between the rendered image and the test scene image, while keeping the Gaussian parameters fixed. This process typically converges within a minute, and is akin to pose estimation for rigid objects but uses motion graphs which generalized to more flexible structures.
Given the reconstructed motion graph of a test scene, we simulate object and the corresponding end-effector movements using object-specific geometric heuristics. For cloth, we simulate diagonal folding by sampling random bending axes and angles; for the microwave, we adjust the relative rotation angle between its two rigid links; and for the rope, we bend its single-chain structure at a randomly selected joint with a sampled angle. Despite these simple heuristics, it is possible to integrate the motion graph simulation with a more powerful physics engine, as a future direction that we discussed in Sec. 5. At each timestep, we evaluate candidate motions by comparing their rendered images against the goal image, then select the motion that maximizes the PSNR. For each task, we evaluate performance across ten different goal images, each selected to be achievable through folding or door-adjustment actions. Our method reliably solve the cloth folding and microwave door manipulation tasks (10/10), while rope bending attains a 70% success rate (7/10). We illustrate and discuss a typical failure case in Fig. 12. Fig. A1 demonstrates our Kuka robot performing all three tasks: cloth folding, microwave door adjustment, and rope bending.
A.4 Ablation Studies
@@
Figure A2: Relationship between novel view rendering quality and motion graph size.
@@
Figure A3: Visualization of rendered images and motion graphs with and without 2D human keypoint regularization (left) and canonical frame regularization (right).
@@
Motion Graph Size. Fig. A2 illustrates the relationship between novel view rendering quality and the number of joints in a motion graph. We observe that the best rendering quality for the teddy scene is achieved with a motion graph of 200 joints. However, increasing the number of joints to 1000 leads to a noticeable drop in quality, suggesting the existence of an optimal graph size. This finding highlights the potential of learning adaptive graph structures that dynamically adjust to different objects.
@@
2D Keypoints Regularization. Fig. A3 compares rendered images and motion graphs with and without 2D human keypoint regularization. We observe that enforcing consistency between the projected 3D kinematic tree and 2D human keypoints (detected by SAPIENS [25]) improves performance, reducing LPIPS by 0.01 to 0.02. More importantly, this regularization produces cleaner and more meaningful motion graphs that align well with the human 3D structure. Without 2D keypoint regularization, the motion graph tends to drift or produce suboptimal structures. For example, if two fingers move together, the kinematic tree may incorrectly control both with a single finger link, leaving other finger links static. This highlights the importance of 2D keypoint regularization in learning accurate and interpretable 3D motion graphs.
@@
Canonical Frame Regularization. The right side of Fig. A3 compares rendered images and motion graphs with and without canonical frame regularization, which minimizes the distance between the learned motion graph’s joint positions and their initial positions at the canonical frame. The joints and links of the motion graph are color-coded based on their distance to the Gaussians on the object surface. With canonical frame regularization, the motion graph remains closely aligned with the object surface. Without it, however, the graph joints become spiky (indicated by red-colored joints) and drift outward, detaching from the object.
Figure A4: Visualization of novel view rendering on iPhone dataset compared with other methods. Regions in green are excluded from evaluation, as these pixels are never seen in training videos. LPIPS on bottom left with best colored red. Our method renders sharper, more complete, and perceptually higher-quality novel views.
Figure A5: Visualization of results in HyperNerf dataset compared with other methods. LPIPS on bottom left with best colored red. Our method achieves comparable rendering quality and LPIPS with state-of-the-art on the chicken, 3D printer, and broom scenes (rows 2 to 4). On the peel-banana scene, our method produces lower-quality results.